benn41:

500 million houses in 42 hours how many per minute

jhonyy9:

like a first step you need convert the 42 hours in minutes do you know it how ?

jhonyy9:

so using the rule of 3 simple you can write - hope you now that 1 hour has 60 minutes 1 hour has 60 minutes 42 hours has x minutes and now cross multiplie so calcule the x value x = 42*60/1 = 42*60 = ? minutes than you get this result you need using again the rule of 3 simple and write 500 million houses in this get result of x minutes 1 house in y minutes so cross multiplie and calcule the y value y = 1*x /500 million = x/500 million minutes hope helped understandably easy